OpenStudy (anonymous):

I will award a medal to anyone who can help with a few questions (:

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

1/3y= 7/3x + 5/3 x-3y=5 3x+y=-13 -2x+5y=-54

OpenStudy (cggurumanjunath):